Convert into a mole.
22 g of Carbon dioxide.
Convert into a mole.
22 g of Carbon dioxide.
 
 

Grade:12th pass

2 Answers

Harshit Singh
2 years ago
Dear Student

Given mass of Carbon dioxide = 22 g
Molar mass of carbon dioxide (CO2) = (1 × 12) + (2 × 16)
= 12 + 32 = .44 g
Mole of carbon dioxide= 22/44 = 0.5 mole

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
Given mass of Carbondioxide = 22 g
Molar mass of Carbondioxide = 44 g
No. Of moles = Given mass / Molar mass
= 22/44 = 1/2 = 0.5 moles
 
Hope it helps!
Thank you ✌️